Hi, I have a problem from my College Algebra homework that I need help with. Logically I think I managed to get the answer to Bob's salary for 2008, being $56K, but I am having a problem with the rest. Thank you for your help.
The following graph shows Bob’s salary from the year 2002 to the year 2005. He was hired in the year 2002; therefore t = 0 represents the year 2002.
The left of the graph is: 40000 - 50000 in 2000 increments. The bottom is 0 - 4 in 1 increment. The line is:0=4100; 1=43,5000; 2=46000; 3=48500.

a) List the coordinates of any two points on the graph in (x, y) form.
(___, ___),(___, ___)
b) Find the slope of this line:
Answer:

Show your work here:
c) Find the equation of this line in slope-intercept form.
Answer:
Show or explain your work here:
d) If Bob’s salary trend continued, what would his salary be in the year 2008?
Answer:
Show or explain your work here:

Answer by scott8148(6628)

a) (1,43500), (3,48500)

b) slope=(change in y)/(change in x) ... m=(48500-43500)/(3-1) ... m=2500
... units are dollars per year ($/yr)

c) slope-intercept is y=mx+b; where m is the slope and b is the y-intercept (value of y when x=0)
... y=2500x+41000
... at zero years (when Bob starts work) his salary is 41000
... his salary increases 2500 each year

d)y=2500(6)+41000 ... y=56000